Annihilator method obtain a particular solution y"'+y=cosx+sinx on R
1
Expert's answer
2021-08-30T17:21:32-0400

Let us write the equation "y'''+y=\cos x+\sin x" in the operator form: "(D^3+1)y=\cos x+\sin x."

The annihilator operator for the right part is "D^2+1." Then we get "(D^2+1)(D^3+1)y=(D^2+1)(\cos x+\sin x)=-\cos x-\sin x+\cos x+\sin x=0."

In the annihilator method for finding a particular solution the method of undetermined coefficients is used. Since "i" and "-i" are not roots of the characteristic equation "k^3+1=0" of the homogeneous differential equation "y'''+y=0," we conclude that the particular solution of the differentail equation "y'''+y=\cos x+\sin x" is of the form "y_p=A\cos x+B\sin x." Then "y_p'=-A\sin x+B\cos x,\ y_p''=-A\cos x-B\sin x,\ y_p'''=A\sin x-B\cos x." Thus we get the equation "A\sin x-B\cos x+A\cos x+B\sin x=\cos x+\sin x," which is equivalet to "(A-B)\cos x+(A+B)\sin x=\cos x+\sin x." Therefore, "A-B=1" and "A+B=1," and hence "A=1" and "B=0."

We conclude that the particular solution is "y_p=\cos x."